OpenStudy (anonymous):

. Multiply. 1 and one-third times 5 and five-eighths

OpenStudy (anonymous):

Answer: = 4/3 × 45/8 = (4 × 45) / (3 × 8) = 180/24 = 15/2 = 7 1/2

OpenStudy (anonymous):

your final answer would be 7 1/2 in simplist form

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

No direct answers please.